OpenStudy (megannicole51):

let u=<4, -1> and v=<3, 5>. what is the absolute value of u+v rounded to the nearest hundredth?

jimthompson5910 (jim_thompson5910):

if u = <a,b> and v = <c,d> then u+v = <a+c, b+d>

jimthompson5910 (jim_thompson5910):

and if z = <p,q> then |z| = sqrt( p^2 + q^2 )

OpenStudy (megannicole51):

the answer choices are a) 3.3 b) 6.1 c) 8.1 d) 9.2

jimthompson5910 (jim_thompson5910):

what did you get for u+v

OpenStudy (megannicole51):

7, 4

jimthompson5910 (jim_thompson5910):

now use the idea that if z = <p,q>, then |z| = sqrt( p^2 + q^2 )

jimthompson5910 (jim_thompson5910):

in this case, p = 7, q = 4

OpenStudy (megannicole51):

49+16 which is 65

jimthompson5910 (jim_thompson5910):

then take the square root of that

OpenStudy (megannicole51):

8.1! awesome thank you so much!

jimthompson5910 (jim_thompson5910):

you got it, yw

OpenStudy (unklerhaukus):

\[\boxed{\huge\color{red}\checkmark}\]